OpenStudy (anonymous):

And for this equation Simplify: (4b + 3) (2b - 5) is the answer 8b2 - 15 ?

OpenStudy (wasiqss):

wrong!!

OpenStudy (anonymous):

no you need 4 multiplications, not two

OpenStudy (anonymous):

\[(4b + 3) (2b - 5) =4b\times 2b-5\times 4b+3\times 2b-3\times 5\]

OpenStudy (anonymous):

so like 8b^2 - 15 ?

OpenStudy (anonymous):

\[8b^2-20b+6b-15\] \[8b^2-14b-15\]

OpenStudy (anonymous):

btw this is not "simplifying" it is "multiplying"

OpenStudy (paxpolaris):

\[\because (4b+3)(2b−5) =\large 4b(2b-5)+3(2b-5)\]